OpenStudy (anonymous):

Find f prime for each: 1.) f(x)=5x ^{6} 2.) f(x) = 8x ^{3} - 10x ^{2} + 7x -5

OpenStudy (anonymous):

\[3.) f(x) = -2/x ^{4} 4.) f(x)= (4x-3)^{2} 5.) f(x)= x-3\sqrt{x} 6.) f(x)= 2x-3/5x+6\]